OpenStudy (anonymous):

HELP!!! Sketch the graph of f based on the following characteristics: - f(0) = f(2) = 0. - f'(x) < 0 if x < 1. - f'(x) = 0. - f(x) > 0 if x > 1. - f''(x) > 0.

OpenStudy (goformit100):

Are there the graphs of in-equality

OpenStudy (anonymous):

i bet it is a parabola with a positive leading coefficient maybe make a quadratic

OpenStudy (loser66):

f(o) =f(2)=0 are x intercepts

OpenStudy (anonymous):

this statement - f'(x) = 0. is a bit confusing i think it is a typo

OpenStudy (anonymous):

i bet it says \(f'(1)=0\)

OpenStudy (anonymous):

Yeah, you are right! It's f'(1) = 0! My mistake! Sorry!

OpenStudy (anonymous):

check \(f(x)=x^2-2x\) and see if it fits all your conditions

OpenStudy (anonymous):

reason being that since the zeros are at \(x=0\) and \(x=2\) you know it looks like \[f(x)=ax(x-2)=ax^2-2ax\] the leading coefficient is positive, since the second derivative is \(2a\) which you are told is positive, and the first coordinate of the vertex is \(1\) which will give you \(a=1\) as well

OpenStudy (anonymous):

YEAH! That's totally right! I was checking f(x) = x^2 - 2x, and it fits all the conditions! Thank you so much! I really appreciate it!

OpenStudy (anonymous):

Also, this is what I got, too: f(x) = ax(x-2)=ax^2 - 2ax.

OpenStudy (anonymous):

I thought it was a parabola, but I wasn't sure; now I am sure!